OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 12 -2x + 3y = 6 -14 24 14 -24

OpenStudy (anonymous):

do you know how to solve these kind of problems?

OpenStudy (anonymous):

not at all

OpenStudy (anonymous):

ok first off try plugging in each # for the y value

OpenStudy (anonymous):

Not 24

OpenStudy (anonymous):

3x-6(-14)=12 then work it out 3x-(-84)=12 (-84)+12=13 3x=13 x=

OpenStudy (anonymous):

x= 4.3333333333

OpenStudy (wolf1728):

A) 3x - 6y = 12 B) -2x + 3y = 6 Multiply equation B by 2 -4x + 6y = 12 and add this to Equation A 3x -6y = 12 -x =24 x = -24 3*-24 - 6y = 12 -6y = 12 +72 -6y = 84 y = -14

OpenStudy (anonymous):

@wolf1728 you shoudlve let her work it out5 or shell never learn

OpenStudy (anonymous):

Thanks